# SCMPPI: Supervised Contrastive Multimodal Framework for Predicting Protein-Protein Interactions

### v1

@@ Abstract
Protein-Protein Interaction (PPI) prediction is a key task in uncovering cellular functional networks and disease mechanisms. However, traditional experimental methods are time-consuming and costly, and existing computational models face challenges in cross-modal feature fusion, robustness, and false-negative suppression. In this paper, we propose a novel supervised contrastive multimodal framework, SCMPPI, for PPI prediction. By integrating protein sequence features (AAC, DPC, CKSAAP-ESMC) with PPI network topology information (Node2Vec graph embedding), and combining an improved supervised contrastive learning strategy, SCMPPI significantly enhances PPI prediction performance. For the PPI task, SCMPPI introduces a negative sample filtering mechanism and modifies the contrastive loss function, effectively optimizing multimodal features. Experiments on eight benchmark datasets, including yeast, human, and H.pylori, show that SCMPPI outperforms existing state-of-the-art methods (such as DF-PPI and TAGPPI) in key metrics such as accuracy ( 98.01%) and AUC (99.62%), and demonstrates strong generalization in cross-species prediction (AUC > 99% on multi-species datasets). Furthermore, SCMPPI has been successfully applied to CD9 networks, the Wnt pathway, and cancer-specific networks, providing a reliable tool for disease target discovery. This framework also offers a new paradigm for multimodal biological information fusion and contrastive learning in collaborative optimization for various combined predictions.
@@ 1 Introduction
To address these issues, this study proposes a novel supervised contrastive multimodal framework (SCMPPI), aimed at further improving PPI prediction performance by combining multimodal features with supervised contrastive learning. We perform sequence embeddings (AAC、DPC、CKSAAP-ESMC) and combine them with the Node2Vec graph embedding method to promote joint representation learning of sequence features and network information. For the PPI task, we have improved the supervised contrastive learning method by adding a negative sample filtering mechanism and modifying the contrastive learning loss function to enhance the model’s ability to embed both sequence and structural information, thus reducing the false-negative rate. Experimental validation on multiple benchmark datasets shows that SCMPPI outperforms existing methods in terms of prediction accuracy, robustness, and generalization ability.
@@ 2 Related Works
In the field of protein-protein interaction (PPI) prediction, the exploration of multimodal models has driven the field toward greater technological complexity and data diversity. By integrating sequence, structural, and network information, these methods significantly enhance prediction accuracy. Models such as DF-PPI (Deep Fusion-PPI) and TAGPPI have gained attention due to their advantages in combining multiple sources of information. TAGPPI (Bosheng Song et al., 2022)[11] innovatively combines AlphaFold-predicted structural features with sequence information to construct a graph neural network model. However, its performance is limited by the accuracy of the structural predictions, and its robustness in low-quality data scenarios still requires further validation. DF-PPI (Hoai-Nhan Tran et al., 2024)[10]enhances sequence feature capture through an improved APAACplus descriptor and integrates handcrafted features with semantic embeddings using a dynamic weighted fusion strategy. It also improves model stability through batch normalization. However, its reliance on GPU resources and the redundancy of handcrafted features may limit its large-scale application.
@@
In this study, we apply supervised contrastive learning [15]to the PPI task. By pulling positive sample pairs closer, pushing negative sample separation farther apart,and extracting label information, we have improved the ability of the multimodal model to embed sequence and structural information. We also incorporate a negative sample filtering mechanism to reduce the likelihood of false negatives and modify the loss of SuoCons $\mathcal{L}^{Sup}_{out}$ to $\mathcal{L}^{P-Sup}$ for PPI.
3 Approach
This section delineates the architecture of the SCMPPI framework, an innovative concept devised to facilitate efficient and high-quality PPI prediction.The structural and design principles of the framework are meticulously expounded in Section 3.1, followed by an overview of the application of multimodal fusion techniques in Section 3.2, and finally its core contrast learning module (Section 3.3) and classifier (Section 3.4), respectively.
@@ 3.1 Architecture of SCMPPI
Figure 1: The Architecture of SCMPPI
@@ 3.2 Multimodal Fusion Encoder
Overall, the encoder extracts protein interaction graph information through Node2vec and protein sequence information through the Sequence Encoder, and then integrates them into a fixed-length protein representation using MLF.
@@
The proposed sequence encoder transforms protein sequences into high-dimensional vector representations for downstream tasks such as function prediction, classification, or structural analysis. It integrates three feature extraction strategies: AAC, DPC, and CKSAAP-ESMC, capturing a range of sequence features, from global composition to local structural patterns and long-range dependencies.
@@
(iii) CKSAAP(k-gap amino acid pairs) [26]uses a set of 400 possible pairings normalized by the rate of k-type amino acid pairs, extending the amino acid pairs to include pairs with a gap of $k$ residues, resulting in a [20, 20, 1] embedding. The frequency is calculated by:
$\text{CKSAAP}(k)=\frac{\text{count of }(aa_{i},aa_{j})\text{ pairs at distance}k}{\text{total pairs at distance }k}$
where $k$ specifies the gap, ranging from $k=0$ to $k=4$. To further extract sequence information, CKSAAP is combined with the ESMC[16] embedding, which extracts the embedding information of amino acid pairs at different intervals. This method iterates over each possible dipeptide pair in the sequence for each gap $k$, computes the average of their embedding vectors, and normalizes according to the gap $k$, ultimately generating a four-dimensional tensor that contains embedding information of dipeptide pairs at various gaps.
Specifically, protein input to the ESMC model produces amino acid-level embeddings $(ae_{1},ae_{2},\dots,ae_{L})$ with a shape of [L, 960]. The mean of amino acid pair embeddings replaces frequency in the formula to produce:
@@
This produces an embedding of size [20, 20, 960], and traversing the distance $k$ from 0 to 3 results in a tensor with a shape of $[k+1,20,20,960]$. This tensor can be reshaped based on downstream tasks, adjusting to a 3D shape $[k\times 960,20,20]$ for 3D convolution operations. In our work, with $k=3$, the final reshaped embeddings have a shape of [3840, 20, 20].
This integration utilizes the biophysical and evolutionary insights of the ESMC model, embedding the contextual information of amino acid pairs up to a gap of $k=3$, reflecting residue correlations and highlighting local interactions within the protein.
In brief,AAC provides the overall amino acid composition, DPC details local structure through adjacent pairs, while CKSAAP with ESMC embeddings captures long-range dependencies and deep contextual features. This fusion across multiple scales—composition, local structure, and global context—ensures a comprehensive and information-rich sequence representation, making it ideal for addressing advanced bioinformatics challenges. For specific parameter choices, refer to Section 3.2.
@@ 3.3 Contrastive Learning Module
Figure 2: Before Contrastive Learning
Figure 3: After Contrastive Learning
@@
We use the Uniform Manifold Approximation and Projection(UMAP)[29] to reduce the dimensionality of the projections, which results in different protein embeddings in a two-dimensional space. In Figure 4, we randomly selected 15 pairs of binding (positive) proteins from the independent test set and projected their corresponding embeddings onto 2D UMAP plots, both before and after contrastive learning. A comparison of the two plots shows that the embeddings are closer together after contrastive learning, demonstrating the effectiveness of contrastive learning in PPI (protein-protein interaction).
@@ 3.4 Prediction Module
4 Experiments
In this section, we describe the experimental setup we used and analyse what we found.
@@ 4.1 Experimental Settings
In our experiments, we represent our algorithm as OURS and compare it with traditional optimizers and existing robust methods for PPI prediction. The traditional optimizers considered are Decision Tree (DT) and Random Forest (RF). The existing robust methods for PPI prediction include DeepFE-PPI[7], PIPR[8], KSGPPI[30], TAGPPI[11], DF-PPI[10], DeepTrio[31], DeeP-AAC[9], DeeP-CNN[9], DFC[32, 33], DCONV[34, 33], and HNSPPI[33]. In this experiment, we utilize a five-fold cross-validation method, which has been widely adopted in previous studies[35, 36, 37]. Additionally, as mentioned earlier, the best configurations of the above models were established as described in their respective works.
@@
Parameter Settings
The model utilizes convolutional layers (two Conv2D layers and a pooling layer) for feature extraction, followed by fully connected layers for sequence representation. It is trained using the AdamW optimizer[44] with a learning rate of 0.001 and a batch size of 32. The loss function comprises binary cross-entropy and contrastive loss, with coefficients set at 0.3, 0.6, or 1. Early stopping is employed with a patience of 3 or 5 epochs, depending on the respective phase of training. The model’s performance is assessed based on the Matthews correlation coefficient (MCC), and the best-performing model is retained. Training is conducted over 30 epochs to ensure the generation of robust sequence representations that facilitate accurate protein-protein interaction prediction.For detailed Parameter Settings , refer to subsection A.2.
Evaluation Configuration
@@ 4.2.1 Performance of SCMPPI on intraspecies dataset
Table 1 shows the 5-fold cross-validation results of the SCMPPI model on the intraspecies datasets. As seen in the table, the SCMPPI model performs excellently and stably across different datasets. The accuracy (acc) and sensitivity (sen) evaluated on the Yeast dataset are both above 98%. AUC and AUPRC represent the model’s ability to distinguish between positive and negative samples at different thresholds and its performance evaluation when handling imbalanced data. The AUC and AUPRC values for SCMPPI on the four intraspecies datasets are all above 90%, demonstrating that our model has high discriminative power and the ability to handle imbalanced data across various datasets, providing a reliable solution for protein-protein interaction prediction.
@@ 4.2.2 Comparisons with existing algorithms
Table 2: Performance comparison on dataset Yeast (clc=0.6).
@@
To validate the superior performance, we compared the proposed model, SCMPPI, with existing robust methods for PPI prediction on all four species-specific benchmark datasets. As shown in Tables 2, 8, 9, and 10, SCMPPI delivers the best performance across the Yeast, Human, H. pylori, and PIPR-cut datasets.
@@ 4.2.3 Robustness and Generalization Study
Test dataset Model AccPreF1% AUCAUPRCMCC
H.pylori OURS 58.41 59.04 58.74 60.04 57.04 0.168
KSGPPI 56.60 58.20 54.30 59.32 56.63 0.134
DeepPE-PPI 48.40 47.71 27.45 48.08 48.74 -0.030
Human OURS 55.26 55.36 54.82 57.23 55.79 0.1052
KSGPPI 53.59 52.41 62.70 55.57 55.01 0.0823
DeepPE-PPI 49.04 46.11 18.30 46.98 47.66 -0.0292
For protein-protein interaction (PPI) prediction, robustness and generalization are crucial because the model needs to accurately predict PPIs across different datasets and data distributions. Therefore, we evaluated the robustness and generalization of our model, and the results show that SCMPPI performs exceptionally well in both aspects.
@@ 4.2.4 Ablation Study
Through ablation experiments, we assessed the contributions of the sequence module, graph module, and contrastive learning module in protein-protein interaction (PPI) prediction. The results from ablation experiments conducted on multiple datasets (Yeast, Human, H. pylori, and PIPR-cut), as shown in Table 12, indicate that the combination of the sequence (Seq), graph (Graph), and contrastive learning (Cl) modules significantly enhances the model’s performance. In all of the datasets that were analysed, the model demonstrated the greatest performance when employing a combination of the three modules (Seq, Graph and Cl). The model demonstrated quasi-optimality in a range of metrics, including accuracy (Acc), the Matthews correlation coefficient (MCC), sensitivity (Sen), the area under the curve (AUC), and the area under the precision-recall curve (AUPRC). When any of the modules (Seq, Graph, or Cl) is removed, the performance drops significantly, highlighting the indispensable contribution of each module to the overall prediction capability.
@@ 4.2.5 Testing on PPI Network Datasets
In this study, we used the SCMPPI model to predict protein-protein interactions (PPIs) in the CD9, Wnt-related pathway, and Cancer-specific networks, successfully predicting all the relevant interactions. Through these predictions, we identified several key proteins that are biologically significant and show potential in clinical applications[45, 46].
Firstly, CD9 is a tetraspanin transmembrane protein widely distributed across various cell types, involved in processes such as cell adhesion, migration, and fusion source[47]. In tumor metastasis, CD9 plays a dual role. In some cancers, such as prostate cancer, high expression of CD9 is associated with lower metastatic ability, exhibiting a tumor-suppressive effect. However, in other types of cancer, such as pancreatic cancer, high CD9 expression correlates with poor prognosis. Therefore, the role of CD9 in tumor prognosis and treatment should be analyzed based on specific cancer types.
The Wnt-related pathway plays a crucial role in embryonic development, cell proliferation, differentiation, migration, and tissue homeostasis. Abnormal activation of this pathway is closely associated with the onset and progression of various cancers source[48]. AXIN1 and WNT9A are core proteins in this pathway. AXIN1 inhibits the activity of the Wnt pathway by forming a complex that degrades β-catenin, while WNT9A, as a ligand protein, activates the pathway. Regulating the activity of the Wnt signaling pathway holds potential clinical value for cancer prevention and treatment.
Finally, in the Cancer-specific network, CDK1 and TP53 are two key proteins. CDK1 regulates cell proliferation during the G2/M phase of the cell cycle, and its abnormal activation can promote tumor cell proliferation source[49]. TP53, as a tumor suppressor gene, is responsible for DNA repair and apoptosis, and mutations in TP53 are commonly found in various cancers source[50]. Studying these proteins and their interactions can provide an important theoretical foundation for cancer therapy.
Through the accurate predictions of the SCMPPI model, we have not only validated the important roles of these core proteins in various biological processes but also opened up new research directions for the early diagnosis and targeted treatment of cancer.
@@ A.2 Parameter settings
A.2.1 The contrastive loss coefs (clc) on model
@@
To evaluate the impact of the contrastive loss coefficient (clc) on model performance, we conducted a grid search on the hyperparameter $\kappa$ in equation (12) across four benchmark datasets (Yeast, Human, H. pylori, and PIPR-cut), with values set as 0, 0.3, 0.6, 1.0. The experimental results, shown in Tables 4567, indicate significant differences in various performance metrics for different settings of $\kappa$. The results demonstrate that the appropriate introduction of contrastive loss (e.g., $\kappa=0.6$ or $\kappa=1.0$) can effectively enhance the model’s generalization ability and robustness, validating the effectiveness of the supervised contrastive learning strategy in multimodal feature fusion.
@@
Table 4: The impact of contrastive loss coefs (clc) on model output (Dataset: Yeast)
Acc Pre Recall Spe F1 MCC AUC AUPRC
clc=0 0.980157 0.978292 0.982123 0.978190 0.980197 0.960336 0.995379 0.995277
clc=0.3 0.980783 0.976969 0.984805 0.976760 0.980858 0.961624 0.996737 0.995502
clc=0.6 0.980068 0.977269 0.983018 0.977118 0.980126 0.960169 0.996201 0.996772
clc=1.0 0.980246 0.977950 0.982659 0.977834 0.980293 0.960517 0.995307 0.994691
Table 5: The impact of contrastive loss coefs (clc) on model output (Dataset: Human)
Acc Pre Recall Spe F1 MCC AUC AUPRC
clc=0 0.828545 0.834707 0.820465 0.836618 0.827163 0.657738 0.898334 0.901011
clc=0.3 0.833931 0.846354 0.816882 0.850980 0.830969 0.66881 0.906778 0.909063
clc=0.6 0.832009 0.829321 0.836621 0.827394 0.832770 0.664347 0.905250 0.908259
clc=1.0 0.830213 0.818392 0.849455 0.810978 0.833354 0.661408 0.905795 0.908416
Table 6: The impact of contrastive loss coefs (clc) on model output (Dataset: H.pylori)
Acc Pre Recall Spe F1 MCC AUC AUPRC
clc=0 0.872133 0.838160 0.929344 0.813380 0.880601 0.750231 0.938190 0.932346
clc=0.3 0.872823 0.838072 0.928640 0.815493 0.880856 0.750217 0.940444 0.932667
clc=0.6 0.872825 0.839097 0.927296 0.816901 0.880875 0.749772 0.934227 0.925803
clc=1.0 0.876650 0.841162 0.932783 0.819014 0.884489 0.757842 0.939882 0.931300
Table 7: The impact of contrastive loss coefs(clc) on model output (Dataset: PIPR-cut)
Acc Pre Recall Spe F1 MCC AUC AUPRC
clc=0 0.883441 0.876867 0.892799 0.874081 0.884588 0.767319 0.939408 0.919204
clc=0.3 0.885669 0.871867 0.904837 0.866508 0.887874 0.772216 0.942730 0.920376
clc=0.6 0.885448 0.881158 0.891236 0.879658 0.886094 0.771080 0.944578 0.924199
clc=1.0 0.883552 0.870030 0.901935 0.865171 0.885650 0.767708 0.940730 0.919098
A.2.2 The k-Spaced (k) and similarity threshold ($\tau$) on model

### v2

@@ Abstract
Protein-protein interaction (PPI) prediction plays a pivotal role in deciphering cellular functions and disease mechanisms. To address the limitations of traditional experimental methods and existing computational approaches in cross-modal feature fusion and false-negative suppression, we propose SCMPPI—a novel supervised contrastive multimodal framework. By effectively integrating sequence-based features (AAC, DPC, ESMC-CKSAAP) with network topology (Node2Vec embeddings) and incorporating an enhanced contrastive learning strategy with negative sample filtering, SCMPPI achieves superior prediction performance. Extensive experiments on eight benchmark datasets demonstrate its state-of-the-art accuracy (98.13%) and AUC (99.69%), along with excellent cross-species generalization (AUC>99%). Successful applications in CD9 networks, Wnt pathway analysis, and cancer-specific networks further highlight its potential for disease target discovery, establishing SCMPPI as a powerful tool for multimodal biological data analysis.
@@ 1 Introduction
To address these issues, this study proposes a novel supervised contrastive multimodal framework (SCMPPI), aimed at further improving PPI prediction performance by combining multimodal features with supervised contrastive learning. We perform sequence embeddings (AAC,DPC,ESMC-CKSAAP) and combine them with the Node2Vec graph embedding method to promote joint representation learning of sequence features and network information.For the PPI task, we enhance supervised contrastive learning by introducing a negative sample filtering mechanism and modifying the contrastive loss function. These improvements strengthen the model’s ability to embed both sequence and network information, thereby reducing the false-negative rate. Experimental validation on multiple benchmark datasets shows that SCMPPI outperforms existing methods in terms of prediction accuracy, robustness, and generalization ability.
@@ 2 Related Works
In the field of PPI prediction, the exploration of multimodal models has driven the field toward greater technological complexity and data diversity. By integrating sequence, structural, and network information, these methods significantly enhance prediction accuracy. Models such as DF-PPI and TAGPPI have gained attention due to their advantages in combining multiple sources of information. TAGPPI (Bosheng Song et al., 2022)[11] innovatively combines AlphaFold-predicted structural features with sequence information to construct a graph neural network model. However, its performance is limited by the accuracy of the structural predictions, and its robustness in low-quality data scenarios still requires further validation. DF-PPI (Hoai-Nhan Tran et al., 2024)[10]enhances sequence feature capture through an improved APAACplus descriptor and integrates handcrafted features with semantic embeddings using a dynamic weighted fusion strategy. It also improves model stability through batch normalization. However, its reliance on GPU resources and the redundancy of handcrafted features may limit its large-scale application.
@@
In this study, we apply supervised contrastive learning [15]to the PPI task. By pulling positive sample pairs closer, pushing negative sample separation farther apart,and extracting label information, we have improved the ability of the multimodal model to embed sequence and structural information. We also incorporate a negative sample filtering mechanism to reduce the likelihood of false negatives and modify the loss of SuoCons $\mathcal{L}^{Sup}_{out}$ to $\mathcal{L}^{P-Sup}$ for PPI. This section delineates the architecture of the SCMPPI framework, an innovative concept devised to facilitate efficient and high-quality PPI prediction.The structural and design principles of the framework are meticulously expounded in Section 3.1, followed by an overview of the application of multimodal fusion techniques in Section 3.2, and finally its core contrast learning module (Section 3.3) and classifier (Section 3.4), respectively.
@@ 2.1 Architecture of SCMPPI
Figure 1: The Architecture of SCMPPI.(a) SCMPPI includes a protein encoder, interaction classifier, and contrastive learning module.(b) Encoder has sequence encoder and Node2vec.(c) Contrastive module projects embeddings into shared space and calculates loss using cosine distance.(d) Sequence encoder uses AAC, DPC, and ESMC-CKSAAP for features.(e) Predicts interaction by outputting a score.
@@ 2.2 Multimodal Fusion Encoder
Overall, the encoder extracts protein interaction graph information through Node2vec and protein sequence information through the Sequence Encoder, and then integrates them into a fixed-length protein representation using FC.
@@
The proposed sequence encoder transforms protein sequences into high-dimensional vector representations for downstream tasks like function prediction, classification, or structural analysis. It integrates three feature extraction strategies: AAC, DPC, and ESMC-CKSAAP, capturing a range of sequence features, from global composition to local structural patterns and long-range dependencies.
@@
(iii) CKSAAP (k-gap amino acid pairs) [26] protein input to the ESMC model produces amino acid-level embeddings $(ae_{1},ae_{2},\dots,ae_{L})$ with a shape of [L, 960]. The mean of amino acid pair embeddings replaces frequency in the CKSAAP formula to produce:
@@
This produces an embedding of size [20, 20, 960], and traversing the distance $k$ from 0 to 3 results in a tensor with a shape of $[k+1,20,20,960]$. This tensor can be reshaped based on downstream tasks, adjusting to a 3D shape $[k\times 960,20,20]$ for 3D convolution operations. In our work, with $k=3$, the final reshaped embeddings have a shape of [3840, 20, 20].For parameter details, refer to subsection A.2.This integration utilizes the biophysical and evolutionary insights of the ESMC model, embedding the contextual information of amino acid pairs up to a gap of $k$, reflecting residue correlations and highlighting local interactions within the protein.
AAC captures the overall amino acid composition, DPC focuses on local structure through adjacent pairs, and CKSAAP with ESMC embeddings captures long-range dependencies and contextual features. This multi-scale fusion provides a rich, comprehensive sequence representation, ideal for bioinformatics tasks.
@@ 2.3 Contrastive Learning Module
We use the Uniform Manifold Approximation and Projection(UMAP)[29] to reduce the dimensionality of the projections, which results in different protein embeddings in a two-dimensional space. In Figure 2, we randomly selected 15 pairs of binding (positive) proteins from the independent test set and projected their corresponding embeddings onto 2D UMAP plots, both before and after contrastive learning. A comparison of the two plots shows that the embeddings are closer together after contrastive learning, demonstrating the effectiveness of contrastive learning in PPI.
@@ 2.5 Experimental Settings
In our experiments, we represent our algorithm as SCMPPI and compare it with traditional optimizers and existing robust methods for PPI prediction. The traditional optimizers considered are Decision Tree (DT) and Random Forest (RF). The existing robust methods for PPI prediction include DeepFE-PPI[7], PIPR[8], KSGPPI[30], TAGPPI[11], DF-PPI[10], DeepTrio[31], DeeP-AAC[9], DeeP-CNN[9], DFC[32, 33], DCONV[34, 33], and HNSPPI[33]. In this experiment, we utilize a five-fold cross-validation method, which has been widely adopted in previous studies[35, 36, 37]. Additionally, as mentioned earlier, the best configurations of the above models were established as described in their respective works.
@@ 2.6.1 Performance of SCMPPI on intraspecies dataset
Table 1 shows the 5-fold cross-validation results of the SCMPPI model on the intraspecies datasets. As seen in the table, the SCMPPI model performs excellently and stably across different datasets. The accuracy (acc) and sensitivity (sen) evaluated on the Yeast dataset are both above 98%. AUC and AUPRC represent the model’s ability to distinguish between positive and negative samples at different thresholds and its performance evaluation when handling imbalanced data. The AUC and AUPRC values for SCMPPI on the four intraspecies datasets are all above 90%, demonstrating that our model has high discriminative power and the ability to handle imbalanced data across various datasets, providing a reliable solution for PPI prediction.
@@ 2.6.2 Comparisons with existing algorithms
Table 2: Performance comparison on dataset Yeast ($\kappa$=1.0).
@@
To validate the superior performance, we compared the proposed model, SCMPPI, with existing robust methods for PPI prediction on all four species-specific benchmark datasets. As shown in Tables 2, 5, 6, and 7, SCMPPI delivers the best performance across the Yeast, Human, H. pylori, and PIPR-cut datasets.
@@ 2.6.3 Robustness and Generalization Study
Test dataset Model AccPreF1% MCC AUCAUPRC%
H.pylori SCMPPI 58.41 59.04 58.74 0.168 60.04 57.04
KSGPPI 56.60 58.20 54.30 0.134 59.32 56.63
DeepPE-PPI 48.40 47.71 27.45 -0.030 48.08 48.74
Human SCMPPI 55.26 55.36 54.82 0.1052 57.23 55.79
KSGPPI 53.59 52.41 62.70 0.0823 55.57 55.01
DeepPE-PPI 49.04 46.11 18.30 -0.0292 46.98 47.66
For PPI prediction, robustness and generalization are crucial because the model needs to accurately predict PPIs across different datasets and data distributions. Therefore, we evaluated the robustness and generalization of our model, and the results show that SCMPPI performs exceptionally well in both aspects.
@@ 2.6.4 Ablation Study
Through ablation experiments, we assessed the contributions of the sequence module, graph module, and contrastive learning module in PPI prediction. The results from ablation experiments conducted on multiple datasets (Yeast, Human, H. pylori, and PIPR-cut), as shown in Table 9, indicate that the combination of the sequence (Seq), graph (Graph), and contrastive learning (Cl) modules significantly enhances the model’s performance. In all of the datasets that were analysed, the model demonstrated the greatest performance when employing a combination of the three modules (Seq, Graph and Cl). The model demonstrated quasi-optimality in a range of metrics, including accuracy (Acc), the Matthews correlation coefficient (MCC), sensitivity (Sen), the area under the curve (AUC), and the area under the precision-recall curve (AUPRC). When any of the modules (Seq, Graph, or Cl) is removed, the performance drops significantly, highlighting the indispensable contribution of each module to the overall prediction capability.
@@ 2.6.5 Testing on PPI Network Datasets
In this study, we applied the SCMPPI model to predict PPIs within three key biological networks: CD9, Wnt-related pathway, and Cancer-specific networks. The model successfully identified critical PPIs, revealing biologically and clinically relevant proteins [44, 45].
CD9, a widely expressed tetraspanin protein, was identified as a central node. Its role in tumor progression is context-dependent—exhibiting tumor-suppressive effects in cancers like prostate cancer, but linked to poor prognosis in others such as pancreatic cancer [46].
In the Wnt-related pathway, SCMPPI highlighted AXIN1 and WNT9A, both central to Wnt signaling regulation. AXIN1 acts to suppress the pathway by promoting $\beta$-catenin degradation, while WNT9A activates it as a ligand. Dysregulation of this pathway is strongly linked to cancer progression [47].
For the Cancer-specific network, CDK1 and TP53 were predicted as key proteins. CDK1 facilitates cell cycle progression, while TP53, a well-known tumor suppressor, controls DNA repair and apoptosis. Mutations in TP53 are commonly implicated in cancer development [48, 49].
Overall, SCMPPI effectively captures meaningful biological insights and provides valuable support for cancer-related research and therapeutic development.
@@ A.2 Parameter settings
A.2.1 Evaluation Configuration
A.2.2 Model Settings
The model utilizes convolutional layers (two Conv2D layers and a pooling layer) for feature extraction, followed by fully connected layers for sequence representation. It is trained using the AdamW optimizer[50] with a learning rate of 0.001 and a batch size of 32. Early stopping is employed with a patience of 5 epochs, depending on the respective phase of training. The model’s performance is assessed based on the Matthews correlation coefficient (MCC), and the best-performing model is retained. Training is conducted over 30 epochs to ensure the generation of robust sequence representations that facilitate accurate protein-protein interaction prediction.
A.2.3 The contrastive loss coefs ($\kappa$) on model
@@
To evaluate the impact of the contrastive loss coefficient ($\kappa$) on model performance, we conducted a grid search on the hyperparameter $\kappa$ in equation (12) across four benchmark datasets (Yeast, Human, H. pylori, and PIPR-cut), with values set as 0, 0.3, 0.6, 1.0. The experimental results, shown in Tables 4, indicate significant differences in various performance metrics for different settings of $\kappa$. The results demonstrate that the appropriate introduction of contrastive loss (such as $\kappa=0.6$ or $\kappa=1.0$) can effectively enhance the model’s generalization ability and robustness, validating the effectiveness of the supervised contrastive learning strategy in multimodal feature fusion.
@@
Table 4: The impact of contrastive loss coefs ($\kappa$) on model output
Dataset $\kappa$ Acc Pre Recall Spe F1 MCC AUC AUPRC
Yeast 0.0 0.9802 0.9783 0.9821 0.9782 0.9802 0.9603 0.9954 0.9953
0.3 0.9808 0.9770 0.9848 0.9768 0.9809 0.9616 0.9967 0.9955
0.6 0.9798 0.9788 0.9809 0.9787 0.9798 0.9596 0.9956 0.9948
1.0 0.9813 0.775 0.9853 0.9772 0.9814 0.9627 0.9969 0.9969
Human 0.0 0.8285 0.8347 0.8205 0.8366 0.8272 0.6577 0.8983 0.9010
0.3 0.8339 0.8464 0.8169 0.8510 0.8310 0.6688 0.9068 0.9091
0.6 0.8320 0.8293 0.8366 0.8274 0.8328 0.6643 0.9053 0.9083
1.0 0.8302 0.8184 0.8495 0.8110 0.8334 0.6614 0.9058 0.9084
H.pylori 0.0 0.8721 0.8382 0.9293 0.8134 0.8806 0.7502 0.9382 0.9323
0.3 0.8728 0.8381 0.9286 0.8155 0.8809 0.7502 0.9404 0.9327
0.6 0.8728 0.8391 0.9273 0.8169 0.8809 0.7498 0.9342 0.9258
1.0 0.8767 0.8412 0.9328 0.8190 0.8845 0.7578 0.9399 0.9313
PIPR-cut 0.0 0.8834 0.8769 0.8928 0.8741 0.8846 0.7673 0.9394 0.9192
0.3 0.8857 0.8719 0.9048 0.8665 0.8879 0.7722 0.9427 0.9204
0.6 0.8854 0.8812 0.8912 0.8797 0.8861 0.7711 0.9446 0.9242
1.0 0.8836 0.8700 0.9019 0.8652 0.8857 0.7677 0.9407 0.9191
A.2.4 The k-Spaced (k) and similarity threshold ($\tau$) on model
In our experiments using the Human dataset, we optimized two hyperparameters k-Spaced (k) and similarity threshold ($\tau$) in Figure66.
Figure 3: ESMC-CKSAAP module. (a) A protein sequence of length $L$ is passed through a pretrained protein model ESMC to obtain an embedding of size $L\times 960$. (b) The amino acid-level protein embeddings obtained are then mapped into CKSAAP according to a one-to-one correspondence rule, resulting in a new sequence embedding.
From Figure 3, different k generates embedding representations of different sizes. where the value of k ranges from 0, 1, 2, 3, 4, and the experimental results show that the model performs best with k=3 after considering the sensitivity, F1 score, and runtime, so we determine k=3 as the final parameter setting (subsection 2.2).
Figure 4: Negative pairs filtering mechanism